# RT-GuIDE: Real-Time Gaussian splatting for Information-Driven Exploration

### v1

@@ IV-B2 Trajectory Generation
$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max},$
$x(0),\theta(0)=x_{0},\theta_{0},\ x(\tau)\in\mathcal{X}_{goal}$
where $J$ is a function that encodes a positive combination of execution time $\tau$ and the control effort $J_{c}(u(\cdot))$.
We use the unicycle model as the robot dynamics. The robot state $x=[p,\theta]\in\mathbb{R}^{2}\times S^{1}$ consists of its position ($p$) and heading ($\theta$). The control inputs $u=[v,\omega]\in\mathbb{R}^{2}$ consist of linear velocity ($v$) and angular velocity $(\omega)$. Inspired by [46], we solve problem 2 by performing a tree search on the motion primitives tree. Motion primitives are generated with fixed control inputs over a time interval $\delta\tau$ with known initial states. Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs. Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization $d\tau$.
We sample a fixed number of points on each motion primitive to conduct collision checks. Since we have an uncertain map, we relax $x\in X_{free}$, to a chance constraint. First, we measure the distance from the discretized positions
$w\sim\mathcal{N}(\kappa,r_{robot}^{2})\in W|\kappa=x_{t}$
along the search path to all Gaussians
$\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\}$
in the map. Our constraint then boils down to the probability of the distance between the test point and the set of Gaussians in the scene
$d_{g}=\inf||W-G||$
being less than 0 i.e. $P(d_{g}<0)\leq\eta$ where $\eta$ is the tolerable risk. When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $\sigma$ of each Gaussian is scaled by a factor of $\lambda_{g}$. The truth value of the test is determined by comparing the distance to the closest Gaussian with $r_{robot}+\lambda_{g}\sigma$, as illustrated in Fig. 2. The rationale for the latter test is as follows. Assuming that the closest obstacle is unique, the probability of collision with some obstacle is well-approximated by the probability of collision with the nearest obstacle. This holds by virtue of the exponential decay of the Gaussian distribution with the distance from the origin. As a result, specifying an upper bound on the probability of collision is approximately the same as specifying the probability of collision with the closest obstacle.
We develop a GPU-accelerated approach for testing all sampled points with all Gaussians from the map at once while growing the search tree. This allows the real-time expansion of the search tree. The cost of each valid motion primitive is defined as
$J(u(\cdot),\tau):=\lambda_{t}\tau+(v^{2}+w^{2})\tau,$
where $\lambda_{t}$ weights the time cost with the control efforts. Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as
$h(p):=||p_{goal}-p||_{2}/v_{max}.$
We use A* to search through the motion primitives tree and keep top $k$ candidate trajectories for the subsequent information maximization. The information of each trajectory is evaluated with state sequence $[x(\delta\tau),\ldots,x(\tau)]$ on the trajectory according to (5). Finally, the trajectory with the highest information $\left(\sum_{t=0}^{\tau}\xi(x(t))\right)$ is selected and executed by the robot.
@@ IV-C Implementation Details
We implemented the proposed method using PyTorch2.4. Most of our parameters of the Gaussian map were set to the default configuration of [14] for the TUM dataset. For simulation experiments, we used 50 mapping iterations. For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence. The mapping module ran at 1Hz and the planning module ran at 3Hz onboard the robot. We set Gaussian frontier size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$.
V Simulation Results & Ablations
@@
Table I: Simulation Results on 120 iTHOR scenes with 20 steps
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$ t/step $\downarrow$
[dB] [m] [s]
Ensemble 12.589 0.483 0.612 1.064 11.094
FisherRF 13.702 0.558 0.549 1.016 5.347
Ours 13.591 0.533 0.576 0.975 0.046
Environments. We conduct simulation experiments with the AI2-THOR simulator [47] on the iTHOR dataset. The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms.
Evaluation metrics & Baselines. For each scene, we generate a test set of images and evaluate the rendered images from each method against the test set. We evaluate the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [48], Learned Perceptual Image Patch Similarity (LPIPS) [49] on the RGB images and Root Mean-Square-Error (RMSE) on the depth images.
@@
We evaluate the proposed information gain metric against two other view selection methods. For the Ensemble baseline, we train an ensemble of $5$ models with the leave-one-out procedure at each step. We compute the patch-wise variance across the rendered images for each sampled viewpoint and ensemble model, and select the viewpoint with the highest variance as the next action. For FisherRF, we follow the original implementation to evaluate viewpoints based on the Fisher information.
Experiments and results. In our simulation experiments, we set the grid size of reachable positions to 0.25m. At each step, we sample viewpoints with different yaw angles around the current position of the agent as possible actions to take. These viewpoints are then evaluated with the corresponding uncertainty metric and the viewpoint with the highest uncertainty is selected as the next action. We ran experiments on all 120 scenes from iTHOR with 20 steps each. We synthesized novel views according to the test set poses and computed the image metrics against the test set images.
The averaged results are presented in Tab. I. Our information gain metric performs comparably with FisherRF and
Figure 3: Comparison of viewpoint selection methods on environment FloorPlan7 of the iTHOR dataset with 50 steps. The average and variance are computed across 5 different seeds. Our proposed method achieves comparable performance to the other methods while exhibiting significantly smaller variance across runs.
@@
outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time. Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint. This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration. We also compute the mean and variance of the evaluation metrics across 50 steps for 5 different seeds, presented in Fig. 3. We note that our method exhibits significantly lower variance than the baselines. The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
VI Real-world Experiments and Benchmark
Figure 4: A qualitative comparison of the novel views rendered using the map generated by our method, onboard and in real-time, to the baseline and the ground truth. The first row is the rendered color images and the second row is the depth images.
Table II: Quantitative Results of Real-world Experiments.
Methods Env. PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$ Traj. Len. [m] Volume Mappedb [%]
Baseline-Voxel Indoor 1.687 123.9 99.94
Baseline-GS 12.79 0.463 0.481 0.548 0.259 123.9 99.94
@@ VI Real-world Experiments and Benchmark
Ours 13.88 0.534 0.442 0.449 0.290 179.1 100.0
Baseline-Voxel Outdoor 0.965 159.9 100.0
Baseline-GS 14.17 0.678 0.406 0.718 0.282 159.9 100.0
@@
Ours 15.66 0.709 0.368 0.513 0.380 225.2 95.53
@@
a Indoor classes: [floor, chair, table, refrigerator, trashcan]. Outdoor classes: [ground, tree, lamppost, rock, signboard, bench].
b Indoor Volume: 360.0m3. Outdoor Volume: 818.6m3.
VI-A Baselines & Experimental Setup
For a meaningful comparison, we implement a frontier-based exploration algorithm presented in [11] as a baseline method. It uses a 3D voxel map to represent the environment, detect frontiers, and generate plans to the nearest frontier. To improve the performance of the baseline method, we incorporated viewpoint sampling at frontier clusters presented in [27] and used A* to provide a reference path to the frontier for the MPL planner.
We deployed our algorithm and the baseline in 2 different real-world environments for comparison on a Clearpath Jackal robot outfitted with AMD Ryzen 5 3600 and RTX 4000 Ada SFF. In addition, it is equipped with an Ouster OS1 LiDAR for state estimation and Intel Realsense D455 camera for mapping. For the baseline method, we set the voxel resolution to 5cm and limit the frontier cluster size to 10cm. For a fair comparison, the baseline method plans on the voxel map but also maintains a Gaussian map to preserve the quality of view synthesis. To generate the groundtruth data for evaluation, we teleoperated the robot to uniformly sample the environment. The Baseline-Voxel method constructs a voxel map and plans with it. The Baseline-GS method constructs a Gaussian map onboard the robot in addition to the voxel map. We set the budget of Ours-Budget to be the trajectory length of the baseline method. We also carried out experiments for our method with a larger budget.
VI-B Evaluation
Qualitative We evaluate the quality of the Gaussian maps that were constructed in real-time onboard the robot by synthesizing images with the test set poses. Samples of synthesized color images and depth images are presented in Fig. 4. Compared to the Gaussian maps constructed by the baseline method, our synthesized depth images capture the depth value closer to the ground truth images while the color images are sharper with more detail. The rendered views from our method at novel camera poses result in better performance of downstream tasks such as semantic segmentation, details are analyzed in the next section.
Metrics. We computed the same image metrics on the RGB and depth images as in the simulation experiments. We rendered 1000 novel views according to the groundtruth poses from the test set and compared them against the corresponding images. We constrained the budget for each method by the distance traveled. To evaluate the coverage of the environment, we computed the volume mapped from the voxel maps built by each method (in post-process for ours).
Results & Discussion. When the budget is set to the trajectory length of the baseline, our method performs comparably with Baseline-GS. When a larger exploration budget is allowed, the frontier baselines terminate early (since they have explored all available frontiers) but our method is able to significantly improve map reconstruction quality. We attribute this to the information gain metric which enables re-visiting of areas in the environment to further improve map quality. We note that Baseline-Voxel is not able to render RGB images since it does not encode color information and also performs poorly in depth rendering due to the discretization of the voxel map. This further motivates the use of a Gaussian map that contains color in addition to geometric information. Our method optimizes for image rendering quality while still achieving mapped volume comparable with the baseline which prioritizes coverage.
@@ VI-B Evaluation
To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also evaluate the rendered images from each method on the task of semantic segmentation. We used Grounded SAM 2 [50, 51, 52, 53, 54] to perform semantic segmentation on the rendered and groundtruth images and computed the mean Intersection over Union (mIoU). Our approach achieves better mIoU scores for both scenes, indicating higher fidelity of the rendered images.
VII Limitations and Future work

### v2

@@ IV-B2 Trajectory planner
$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max}$
where $\lambda_{t}$ weights the time cost with the control effort, $v_{max}$ and $\omega_{max}$ are actuation constraints, and $x_{0}$ is the initial state.
Motion Primitive Tree Generation. Inspired by [45], we solve problem 1 by performing a tree search on the motion primitives tree. Motion primitives are generated with fixed control inputs over a time interval with known initial states. Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs. Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization. Collision Checking. We sample a fixed number of points on each motion primitive to conduct collision checks. Since we have an uncertain map, we relax $x\in X_{free}$ to a chance constraint. Let $d_{\kappa}$ be the minimum distance between a test point and the set of Gaussians. Our constraint then amounts to the probability of the distance between the test point and the set of Gaussians in the scene being less than some allowed tolerance $\gamma$ with probability $\eta$ i.e. $P(d_{\kappa}<\gamma)\leq\eta$. When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $r$ of each Gaussian is scaled by a factor of $\lambda_{g}$. The truth value of the test is determined by comparing the distance to all Gaussians with $\gamma:=r_{robot}+\lambda_{g}r$, as illustrated in Fig. 2. Note that setting $\lambda_{g}=3$ is equivalent to the test proposed by [37]. However, unlike the continuous trajectory optimization approach in [37], which requires subsampling to provide soft constraints, our search-based method allows dense checks. We are implicitly assuming that 3DGS accurately captures geometry of the scene and we notice this holds empirically since we initialize Gaussians in the map using the depth measurements from the RGB-D sensor. In the case of collision checking against non-isotropic Gaussians, the principal axis of the Gaussian can be taken as the radius. We develop a GPU-accelerated approach for testing collisions of all sampled points with all Gaussians from the map at once while growing the search tree. This allows the real-time expansion of the search tree.
Tree Search. The cost of each valid motion primitive is defined by Prob. 1. Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as $h(p):=||p_{goal}-p||_{2}/v_{max}$. We use A* to search through the motion primitives tree and keep top $N_{traj}$ candidate trajectories for the information gain maximization.
Information Gain Maximization. A state sequence $\{x_{i}({l\tau/L})\ |\ l\in\{0,\ldots,L\}\}$ containing the end state of each of $L$ segments in the $i$-th trajectory is used for evaluating the information along the trajectory. The information of each state is evaluated according to eq. (9). Finally, the trajectory with the highest information
$\mathop{\arg\max}\limits_{i\in\{1,\ldots,N_{traj}\}}\sum_{l=0}^{L}\xi\left(x_{i}\left(\frac{l\tau}{L}\right)\right),\ l\in\{0,\ldots,L\}$ (11)
is selected and executed by the robot.
@@ IV-C Implementation Details
We implemented the proposed method using PyTorch2.4 and the mapping module following [14]. We note that our framework is compatible with other Gaussian splatting approaches if they can meet the compute and latency requirements. The parameters of the mapper were set to the default configuration of [14] for the TUM dataset. For simulation experiments, we used 50 mapping iterations. For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence. The mapping and planning modules are both set to 1Hz onboard the robot. For trajectory viewpoint evaluation, we set the threshold for high and low uncertainty Gaussians at half a standard deviation above and below the average magnitude of parameter updates, respectively. We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor. This necessitated the removal of Gaussians on the ground plane for planning. We set the Gaussian region size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$. For planning, we set $N_{v}=3$, $v_{max}=0.6$, $N_{\omega}=5$, $\omega_{max}=0.9$, and time discretization 1s.
V Results
V-A Simulation Experiments & Ablations
@@
Table I: Benchmark experiments on iTHOR test scenes with 100 steps
Methods PSNR SSIM LPIPS RMSE t/step
[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [s] $\downarrow$
Ensemble 21.489 0.762 0.360 0.350 2.667
FisherRF 24.171 0.849 0.270 0.285 1.299
RTGuIDE 22.946 0.820 0.305 0.343 0.013
Table II: Ablation experiments on iTHOR test scenes with 100 steps
Methods PSNR SSIM LPIPS RMSE t/step
[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [s] $\downarrow$
RTGuIDE 22.946 0.820 0.305 0.343 0.013
RTGuIDE-sum 20.279 0.740 0.371 0.512 0.007
RTGuIDE-sq 22.489 0.802 0.308 0.343 0.014
RTGuIDE w. noise 22.541 0.801 0.329 0.344 0.013
RTGuIDE-sq w. noise 21.445 0.779 0.351 0.386 0.013
V-A1 Experiment setup & Baselines
To evaluate our proposed heuristic, we conduct simulation experiments with the AI2-THOR simulator [46] on the iTHOR dataset. The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms. Each step in the simulator involves a translation and rotation. The agent is allowed to translate 0.5m in both $x$ and $y$ positions. At each allowed position, we sample a set of candidate yaw angles to form a discrete set of possible orientations. These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints. Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action. We ran experiments on all test scenes from iTHOR with 100 steps each. The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
@@ V-A1 Experiment setup & Baselines
For each scene, we synthesized novel views from the generated maps from a set of uniformly sampled test poses and computed the image metrics against the ground truth images. We evaluate the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [47], Learned Perceptual Image Patch Similarity (LPIPS) [48] on the RGB images and Root Mean-Square-Error (RMSE) on the depth images. We evaluate the proposed method (denoted as Ours) on viewpoints following eq. (9) against two other view selection methods. For FisherRF, we follow the original implementation to evaluate viewpoints based on the Fisher information. For the Ensemble baseline, we train an ensemble of $5$ models with the leave-one-out procedure at each step. We compute the patch-wise variance across the rendered images for each sampled viewpoint and ensemble model, and select the viewpoint with the highest variance as the next action.
V-A2 Benchmark experiments
@@
The averaged results over all test scenes are presented in Tab. I. Our heuristic performs comparably with FisherRF and outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time. Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint. This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration.
V-A3 Ablation experiments
For ablation experiments, we implemented RTGuIDE-sum, which solely considers the sum of the uncertainty of Gaussians in the camera view. RTGuIDE-sq evaluates the uncertainty of Gaussians as the squared L2 norm following the original derivation in eq. (8) and using eq. (9) to evaluate the uncertainty of viewpoints. To simulate real-world noisy measurements, we added Gaussian noise $(\mu=0,\sigma^{2}=0.1)$ to depth measurements in the simulation to evaluate RTGuIDE w. noise and RTGuIDE-sq w. noise.
As shown in Tab. II, the ablation experiments on different metrics demonstrate that our proposed approach outperforms the simple uncertainty summation (RTGuIDE-sum), which focuses purely on exploitation. Furthermore, the results show that both the squared (RTGuIDE-sq) and standard L2 norms (RTGuIDE) achieve comparable performance when perfect depth measurements are available. However, when noise is introduced, the experiments confirm that our chosen metric is more robust and better suited for real-world scenarios where measurement noise is inevitable.
V-B Planner Performance
We evaluate the necessity of our proposed GPU-based planning approach in enabling real-time planning. In particular, we conducted experiments on trajectory planning with a 5-meter horizon, evaluating both GPU-accelerated collision checking and a fully CPU-based implementation. As illustrated in Fig. 3, the planning time of CPU-based planner increases significantly with the number of Gaussians. For example, a 40m $\times$ 40m outdoor parking lot can contain around $3\times 10^{6}$ Gaussians. In this case, the CPU-based planner requires 11.65 seconds, whereas the GPU-based planner takes 0.62 seconds, achieving 18$\times$ speedup. The GPU-based collision testing enables parallel growth of the search tree at each layer, along with simultaneous and efficient collision checks between all test points and Gaussians directly on the GPU.
@@ V-B Planner Performance
Figure 3: Time spent to plan a trajectory with a 5m horizon (including 129 collision checking points) versus the number of Gaussians in the map.
V-C Real-world Experiments & Benchmarks
Table III: Quantitative Results of Real-world Experiments.
Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$
Ensemble Indoor 1 72m2 300 6.27 0.204 0.839 2.051 0.071
FisherRF 16.80 0.665 0.393 0.242 0.314
RTGuIDE 17.83 0.737 0.334 0.202 0.338
@@ V-C Real-world Experiments & Benchmarks
Ensemble Indoor 2 189m2 600 10.41 0.269 0.787 2.410 0.156
FisherRF 14.88 0.588 0.451 1.170 0.248
RTGuIDE 16.40 0.711 0.299 0.689 0.391
[…]
Ensemble Outdoor 1 252m2 600 10.59 0.392 0.637 1.705 0.549
FisherRF 17.58 0.687 0.349 0.519 0.600
RTGuIDE 20.00 0.744 0.296 0.369 0.622
[…]
Ensemble Outdoor 2 480m2 900 16.79 0.538 0.576 1.422 0.159
FisherRF 19.44 0.680 0.439 0.773 0.177
RTGuIDE 20.27 0.750 0.345 0.399 0.174
[…]
@@
a Indoor 1 classes: [floor, chair, table, refrigerator, cabinet, backpack, plants]. Indoor 2 classes: [floor, chair, table, couch, cushion, trashcan, television, plants]. Outdoor 1 classes: [pavement, plants, barrel, traffic cone]. Outdoor 2 classes: [pavement, grass, tree, bench, lamppost].
V-C1 Experiment setup & Baselines
To evaluate the effectiveness of our proposed framework, we compared it with two baselines onboard the robot. We first implemented FisherRF [36] by constructing both a voxel map and a Gaussian splatting map online. We detect and cluster frontiers from the voxel map and use A* to generate reference paths to frontier clusters. Camera views along each path are evaluated to compute information gain. To generate collision-free and dynamically feasible trajectories, we generate trajectories along waypoints in the path through MPL [45] in the voxel map considering the unicycle model for robot dynamics. We also implemented an Ensemble method and deployed it on the robot, similar to the framework of FisherRF. In the ensemble approach, uncertainty is computed as described in Sec. V-A1.
We deployed our method and the baselines in four different real-world environments on a Clearpath Jackal robot outfitted with an AMD Ryzen 5 3600 and RTX 4000 Ada SFF. In addition, the platform is equipped with an Ouster OS1 LiDAR for state estimation and a ZED 2i stereo camera for mapping. We used [49] to provide odometry and truncated the depth measurements at 5 meters for all methods in all of our experiments. For the baseline methods, we set the voxel resolution to 5cm. To generate the ground truth data for evaluation, we teleoperated the robot to uniformly sample the environment. The exploration budget for all methods is set to $\sim$5x the time needed to fully survey the environment with teleoperation. We rendered 500 novel views according to the ground truth poses and compared them against the corresponding images.
@@ V-C1 Experiment setup & Baselines
We computed the same image metrics on the rendered RGB and depth images as in the simulation experiments. To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also performed an evaluation on the task of semantic segmentation. We used Grounded SAM 2 [50] to obtain segmentation masks of the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
V-C2 Qualitative results
Figure 4: Visualization of onboard constructed map
Figure 5: Qualitative results
We evaluated the quality of the Gaussian splatting maps that were constructed in real-time onboard the robot by rendering novel views with a set of test poses. A visualization of the onboard constructed map is shown in Fig. 4, and examples of rendered color images are presented in Fig. 5. Constructed in real time onboard the robot, our Gaussian splatting map provides a detailed representation of the environment. Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in detail in the next section.
V-C3 Quantitative results
As shown in Tab. III, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot. To validate that our approach generates maps of reasonable quality, we also present maps generated with the ground truth samples. Results on novel view image rendering show that the maps constructed with our approach achieve 0.8-2.4 higher PSNR than FisherRF and 3.5-11.6 higher PSNR than Ensemble. Furthermore, when evaluated on depth image recovery, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble. These results reflect that our method outperforms the baselines in achieving good coverage of the environments while also optimizing for map quality during exploration.
For semantic segmentation, our approach achieves better mIoU scores for most experiments, indicating higher fidelity of the rendered images compared to the baselines.
V-C4 Discussion
We attribute the performance of our approach to two key factors. First, the use of an information metric that is efficient to compute onboard the robot enables smooth and continuous operation. Second, since our method does not rely on frontiers extracted from the voxel map for geometric coverage, it allows for revisiting of areas in the environment to further improve map quality.
VI Limitations and Future work